OpenStudy (anonymous):

can someone help me with these two questions please! Jordan wants to solve the following system using the elimination method: 2x + 3y = 10 x + y = 7 What number should the equation x + y = 7 be multiplied by to eliminate y? −2 2 −3 3 Question 3 (Multiple Choice Worth 4 points) (08.05) A student is trying to solve the system of two equations given below: Equation P: a + b = 6 Equation Q: 4a + 2b = 19 Which of the following steps can be used to eliminate the a term? −1(4a + 2b = 19) −4(4a + 2b = 19) −4(a + b = 6) 4(a + b = 6)

OpenStudy (anonymous):

@Krane103 can you help me please

OpenStudy (anonymous):

@linda360 can you hlp

OpenStudy (anonymous):

@senpainoticedyou hlp please

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

or 3, depends on if you add or subtract

OpenStudy (anonymous):

okay thanks can you help with the other one

OpenStudy (anonymous):

which one do you think it is?

OpenStudy (anonymous):

−4(4a + 2b = 19) mabey

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

best guess

OpenStudy (anonymous):

i dont know how to do this

OpenStudy (anonymous):

are you still there

OpenStudy (anonymous):

The point of studying is understanding, not guessing. The goal of the elimination method is to get rid (elimination is another word for 'getting rid of) of one of the variables. (usually x and y, but it can be a or b or whatever too). For example if we have : 2x + 3y = 10 x + y = 7 we can multiply the second one with -3 so that we have +3y in the top equation, and -3y in the bottom equation -3(x+y=7) = -3x-3y= -21 2x + 3y = 10 -3x-3y= -21 + ------------------------------ -x - 0y = -11 x = 11 x+y = 7 11-7=y y = 4

OpenStudy (anonymous):

There's multiple numbers that work, it doesn't really matter which one you use, as long as you make it easy to either subtract or add the resulting equations in such a way that one of the unknown variables drops out of the equation.